a rocket is divided into three sections. the payload and navigation section in the top, the fuel section in the middle and the rocket engine section in the bottom. the top section is one-sixth the lenght of the bottom section. The middle section is one-half the length of the bottom section. The total length is 240ft. find the lenght of each section
Answer by mananth(16946)

let the bottom section be x
the top will be x/6
the middle will be x/2
total length = 240
..
x+ x/6 + x/2 =240
LCM for left side = 6
6x+x+3x /6 = 240
10x= 1440
x= 144 feet
..
top = x/6 = 144/6
=24 feet
..
middle = x/2 = 144/2 = 72 feet
..
72+24+144= 240